OpenStudy (anonymous):

iron ball bearing with circumference of 3.25mm has mass of 4.20g. calculate density of iron in g/cm3?

OpenStudy (anonymous):

well circumference is volume and density is mass over volume so you just need to recalculate volume in mm3 to cm3 and divide those two numbers... rho = m/V

OpenStudy (jfraser):

@Kryten, circumference isn't volume, but we can use radius to find volume of a sphere\[V = \frac{4}{3}\pi(r^3)\] then plug into the equation from above

OpenStudy (anonymous):

@JFraser then google translate didnt translate it correctly, i apologize for wrong answer...